# Index Terms: Monte-Carlo Tree Search, Proof-Number Search, MCTS SolverI Introduction

### v3

@@ VII PN-MCTS Experiments
To encode which PN-MCTS enhancement is turned on, we use the following notation. Final move selection is encoded by F, Solver by S, and UCT-PN formula by U. If an enhancement is off, the x symbol is used instead. Thus, xSx encodes a variant where only PN-solver is used, while FSU is a variant with all proposed enhancements enabled.
@@ VII-A PN-MCTS Variants
The results of the experiment are presented in Table II. To make the comparison meaningful, in the xxx variant, none of the enhancements are used, but all data required to use them is computed. Thus, this is the baseline, in which the winrates are strictly correlated with the results of the overhead experiment.
@@
Second is that, indeed, proposed enhancements are mostly improvements, and turning them on increases the winrates. It is not always the case, but the FSU variant with all enhancements is the best combination in 3 out of 5 tested domains and second-best in one more.
@@ VII-A1 Final Move Selection
Although the benefits of this enhancement are apparent, and it has always been a crucial part of MCTS-Solver [16], its impact was never measured alone, apart from the selection part of the solver. Our experiments measure this impact, as we can directly compare combinations of enhancements with, and without final move selection: xxx and Fxx, xSx and FSx, xxU and FxU, finally xSU and FSU.
@@ VII-A2 Solving Subtrees
This enhancement is usually beneficial, and quite impactful, providing improvements ranging from about 3 percent points with all other enhancements, to nearly 20 percentage points for LOA and simpler variants of the algorithm, e.g., Fxx to FSx.
@@
However, for Awari, solver tends to be harmful. The highest negative impact observed, over is for the most complex variant, between FxU to FSU. Again, the reason here is the high frequency of draws in the game. Most likely, the impact of excluding of some child nodes from contributing to the UCB1 selection formula.
@@ VII-A3 UCT-PN
In 14 out of 20 comparable variants, winrates increased by more than 10%. The biggest observed improvement, 44.4%, is for LOA , from Fxx to FxU.
@@ VII-B Time Trends
To assess the usefulness of the proposed MCTS variant it is necessary to measure the effect of time and estimate the trends. We have selected all variants with at least two enhancements applied (FxU, FSx, FSU, xSU), and run them for different time limits: 0.125, 0.25, 0.5, 1, 2 and 4 seconds per move. The results of this experiment are presented in Figure 4.
@@
For LOA, independently of the board size, there is generally an improving tendency for variants with the UCT-PN enhancement, and worsening for FSx.
The peak winrate with 96.2% is reached by FSU for the larger LOA board and the 2 seconds time limit.
For MiniShogi variants with final move selection perform roughly stable for all times, while xSU has poor results for low time limits, and becomes comparable to other variants with s.
The trends for Knightthrough are harder to generalize, with FSx and FSU behaving in U-like shape, and FxU being clearly worse, staying below 60% for all times.

### v4

@@ VII PN-MCTS Experiments
To encode which PN-MCTS enhancement is turned on, we use the following notation. Final move selection is encoded by F
, Solver by S
, and UCT-PN formula by U
. If an enhancement is off, the x
symbol is used instead. Thus, xSx
encodes a variant where only PN-solver is used, while FSU
is a variant with all proposed enhancements enabled.
@@ VII-A PN-MCTS Variants
The results of the experiment are presented in Table II. To make the comparison meaningful, in the xxx
variant, none of the enhancements are used, but all data required to use them is computed. Thus, this is the baseline, in which the winrates are strictly correlated with the results of the overhead experiment.
@@
Second is that, indeed, proposed enhancements are mostly improvements, and turning them on increases the winrates. It is not always the case, but the FSU
variant with all enhancements is the best combination in 3 out of 5 tested domains and second-best in one more.
@@ VII-A1 Final Move Selection
Although the benefits of this enhancement are apparent, and it has always been a crucial part of MCTS-Solver [16], its impact was never measured alone, apart from the selection part of the solver. Our experiments measure this impact, as we can directly compare combinations of enhancements with, and without final move selection: xxx
and Fxx
, xSx
and FSx
, xxU
and FxU
, finally xSU
and FSU
@@ VII-A2 Solving Subtrees
This enhancement is usually beneficial, and quite impactful, providing improvements ranging from about 3 percent points with all other enhancements, to nearly 20 percentage points for LOA and simpler variants of the algorithm, e.g., Fxx
to FSx
@@
However, for Awari, solver tends to be harmful. The highest negative impact observed, over is for the most complex variant, between FxU
to FSU
. Again, the reason here is the high frequency of draws in the game. Most likely, the impact of excluding of some child nodes from contributing to the UCB1 selection formula.
@@ VII-A3 UCT-PN
In 14 out of 20 comparable variants, winrates increased by more than 10%. The biggest observed improvement, 44.4%, is for LOA , from Fxx
to FxU
@@ VII-B Time Trends
To assess the usefulness of the proposed MCTS variant it is necessary to measure the effect of time and estimate the trends. We have selected all variants with at least two enhancements applied (FxU
, FSx
, FSU
, xSU
), and run them for different time limits: 0.125, 0.25, 0.5, 1, 2 and 4 seconds per move. The results of this experiment are presented in Figure 4.
@@
For LOA, independently of the board size, there is generally an improving tendency for variants with the UCT-PN enhancement, and worsening for FSx
The peak winrate with 96.2% is reached by FSU
for the larger LOA board and the 2 seconds time limit.
For MiniShogi variants with final move selection perform roughly stable for all times, while xSU
has poor results for low time limits, and becomes comparable to other variants with s.
The trends for Knightthrough are harder to generalize, with FSx
and FSU
behaving in U-like shape, and FxU
being clearly worse, staying below 60% for all times.